You take out a loan of $15,000 to purchase a car.
If your loan is for 3 years at a simple interest rate of 10.5%, find out how much interest and the total amount paid for the loan.

Found 2 solutions by Boreal, MathTherapy:
Answer by Boreal(15235)

I=prt
I=15000*0.105*3=$4725 interest paid
total amount paid back is that plus the original loan or $19,725.

Answer by MathTherapy(10552)

You take out a loan of $15,000 to purchase a car.
If your loan is for 3 years at a simple interest rate of 10.5%, find out how much interest and the total amount paid for the loan.
Total amount paid, or Future Value of the $14,000 loan, or FV = P(1 + TR) = 14,000(1 + 3 * .105) = $19,725.

Of the $19,725 paid, $15,000 was the principal or loan amount, so interest paid on the loan = $19,725 - $15,000 = $4,725.